HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

EXPERIMENTAL VALIDATION OF CO_{2} LINE LISTS IN THE SPECTRAL RANGE NEAR 0.69 μm

\underline{K. BIELSKA} , A. A. BALASHOV, S. WÓJTEWICZ, R. CIURY� O, D. LISAK, Institute of Physics, Faculty of Physics, Astronomy and Informatics, Nicolaus Copernicus University in Toruń, Grudziadzka 5, 87-100 Torun, Poland; A. RAJ, Department of Applied Chemistry and Institute of Molecular Science, National Yang

Ming Chiao Tung University, Hsinchu, Taiwan 30010

We conducted an attempt to measure CO_{2} lines in the spectral range between 0.687 and 0.699 μm. The predicted line intensities are below 2.1×10^{−30} cm/molecule. To detect these weak lines we employed a cavity ring-down spectroscopy (CRDS) technique. It was already shown that our spectrometer has sufficient sensitivity to measure a CO line having intensity of 1.807 × 10^{−30} cm/molecule with a signal-to-noise ratio of 36 if very high-reflectivity ring-down cavity mirrors are used^{1}.

CO_{2} lines having intensities of up to 2.076×10^{−30} cm/molecule are predicted by the ExoMol^{2} line list and available in the HITRAN 2020^{3} database. Also, the Ames-2021 line list^{4} contains lines located in this spectral range having intensities up to 1.819×10^{−31} cm/molecule. We probed altogether five spectral ranges corresponding to relatively strong lines among ones predicted by both line lists. Within these ranges, there are expected more than twenty CO_{2} lines. Our estimated detection limit in the case of each probed spectral range is between 2 and above 10 times lower than the expected peak absorption level. However, none of these lines was detected in our experiment. Our result is in contrast to recent FTS measurements^{5} for two orders of magnitude stronger CO_{2} bands at slightly higher wavelengths. However, in the case of the HITRAN database, lines detected with the FTS are predicted by a different set of calculations than the ones that we intended to measure.

1A.A. Balashov et al., Measurement and calculation of CO (7–0) overtone line intensities, accepted in J. Chem. Phys.

2doi:10.1093/mnras/staa1874, S.N. Yurchenko et al., ExoMol line lists - XXXIX. Ro-vibrational molecular line list for CO_{2}, Mon. Not. R. Astron. Soc. 496, 5282-5291 (2020).

3doi:10.1016/j.jqsrt.2021.107949, I.E. Gordon et al., The HITRAN2020 molecular spectroscopic database, J. Quant. Spectrosc. Radiat. T. 277, 107949 (2022).

4doi:10.1021/acs.jpca.2c01291, X. Huang et al., Ames-2021 CO_{2} dipole moment surface and IR line lists: toward 0.1% uncertainty for CO_{2} IR intensities, J. Phys. Chem. A 126, 5940-5964 (2022)

5doi:10.1016/j.jqsrt.2023.108595, O.M. Lyulin et al., The absorption bands of ^{12}C^{16}O_{2} near 718 nm, J. Quant. Spectrosc. Radiat. T. 303, 108595 (2023).

p-number: p044  Submitted on Wed, 07 Jun 2023 13:10:01 +0200